3. At 250C, the volume of chlorine gas is 15 dm3 .What is the resulting volume if the temperature is adjusted to 320K provided that the pressure remains constant?



1
Expert's answer
2022-05-17T04:06:04-0400

Given:

V1 = 15 dm3

T1 = 25°C + 273 = 298 K

P1 = P2 = constant

V2 = unknown

T2 = 320 K


Formula: V1/T1 = V2/T2


Solution:

Since the pressure and amount of gas remain unchanged, Charles's law can be used.

Charles's law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the resulting volume, solve the equation for V2:

V2 = V1T2 / T1

V2 = (15 dm3 × 320 K) / (298 K) = 16.11 dm3

V2 = 16.11 dm3


Answer: The resulting volume of chlorine gas is 16.11 dm3